OpenStudy (anonymous):

FAN AND MEDAL!!!! m + 3 = 2m. m = 3. Can you please explain how m = 3? Why is 2m - m = 3? Please respond ASAP!!!

pooja195 (pooja195):

m=3 is what your supposed to plug in your equation

OpenStudy (anonymous):

mhmmmm

OpenStudy (anonymous):

i know but how does m = 3?

OpenStudy (anonymous):

i mean 2m - m = 3

pooja195 (pooja195):

i was thinking of pluging in 0 for the 1st m but u would be left with -3......so im not sure

OpenStudy (anonymous):

3 + 3 = 2 x 3 ...... 6 = 6 that makes sense

OpenStudy (anonymous):

but how do i get x = 3

OpenStudy (anonymous):

that is the question

pooja195 (pooja195):

if its that simple then why not plug in 3 into your equation?

pooja195 (pooja195):

This question makes no sense

OpenStudy (anonymous):

i dont understand what you mean by "plug"

OpenStudy (anonymous):

u mean substitute 3 for m?

pooja195 (pooja195):

right same thing :P

OpenStudy (anonymous):

ik like i said m + 3 = 2m, so m has to be one number that makes both equations true... if you get m = 3, you substitute m for 3...

OpenStudy (anonymous):

3 + 3 = 2 x 3

OpenStudy (anonymous):

6 = 6

OpenStudy (anonymous):

THAT i already know, but how do i mathematically get m = 3? on yahoo it said they got 3 cuz 2x - x = 3... they "isolated the variable"

pooja195 (pooja195):

AH!!! im so confused >_< i dont get this problem @tkhunny can you please help :)

OpenStudy (anonymous):

yes please @tkhunny ?//

OpenStudy (anonymous):

so basically you have the answer you just need help understanding why that is the answer?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

WOW should've put it in that sense so does the question ask tell wether the given coordinate is the solution?

OpenStudy (sleepyjess):

First, what is 2m - m?

OpenStudy (anonymous):

they say it's 3 but i dont get why

OpenStudy (anonymous):

i know the answer but i gust dont understand how to get it

OpenStudy (anonymous):

just*

OpenStudy (sleepyjess):

Ok, I don't need to know the value of m, what is 2m - m?

Nnesha (nnesha):

 you need to combine m so that's y move m to the right side so then you can combine with 2m